OpenStudy (anonymous):

Estimate the volume of the solid that lies below the surface z=xy and above the rectangle R. Use a Riemann sum with m=3, n=2, and take the sample point to be the upper right corner of each square. R = {(x,y): 0<=x<=6, 0<=y<=4}

OpenStudy (nathan917):

For The whole question here is the answer. Δx = (6 - 0)/3 = 2 and Δy = (12 - 8)/2 = 2. So, our region R is partitioned as follows: (0, 12) (2, 12) (4, 12) (6, 12) , (0, 10) (2, 10) (4, 10) (6, 10) , (0, 8) (2, 8) (4, 8) (6, 8) --------------- (a) Using upper right endpoints for each of the little 'grids' yields (with f(x,y) = xy) ∫∫R xy dA ≈ 2 * 2 [f(2, 10) + f(4, 10) + f(6, 10) + f(2, 12) + f(4, 12) + f(6, 12)] = 4 [20 + 40 + 60 + 24 + 48 + 72] = 1056. (b) Using the midpoint for each of the little 'grids' yields ∫∫R xy dA ≈ 2 * 2 [f(1, 9) + f(3, 9) + f(5, 9) + f(1, 11) + f(3, 11) + f(5, 11)] = 4 [9 + 27 + 45 + 11 + 33 + 55] = 720.

OpenStudy (anonymous):

@nathan917 thanks for the reply! Would you care to help me through this step by step?

OpenStudy (anonymous):

Firstly, the [\Delta\] x and y are the increments in which our square will be partitioned along the x and y axes?

OpenStudy (anonymous):

I'll just answer my own question then, lol. @nathan917 your answer was actually incorrect, however, it helped me get on course with what is asked. Okay so the m and n is the intervals. So: \[\Delta x=\frac{6-0}{3}=2\]\[\Delta y=\frac{4-0}{2}\] Thus, the /real/ sample points will be: (2,2) (2,4) (4,2) (4,4) (6,2) (6,4) Now \[M(m,n)=M(3,2)=\Delta x \Delta y \sum_{i=1}^{3} \sum_{j=1}^{2}f(x_{i},y_{j})\]\[=4[f(2,2)+f(2,4)+f(4,2)+f(4,4)+f(6,2)+f(6,4)]=288\]